OpenStudy (anonymous):

Determine graphically the x value when y=5 in the given equation. -x+2y=8

OpenStudy (anonymous):

You have fist to draw the representation of the curve that its equation is -x+2y=8. What would it like to be ?

OpenStudy (anonymous):

I dont know what you mean by that

OpenStudy (anonymous):

-x+2y=8 is an equation for a straight line and we can trace it by finding two points from the line and then we draw the line passing by these points. For example : when x=0 we have 0+2y=8, and then y=4. So the point A(0,4) belongs to the line. when y=0 we have -x+0=8, and then x=-8. So the point B(-8,0) belongs to the line. So we draw the line passing by the points A and B. Did you get it ?

OpenStudy (anonymous):

yeah i think so

OpenStudy (anonymous):

Now, when you draw that line, we draw a "horizontal line" passing by the point (0,5) - that is the line which its equation is y=5 - then we find the intersection of the two lines, and then we can find the correspond value of x of the point of the intersection.